# Ground State Energy of Dilute Fermi Gases in 1D

### v1

@@ 4.3.4 Proof of Lemmas 19, 20 and 21
where we used $|W_{12}(x_{1}-x_{2})\chi|^{2}\leq b^{2}$ (see Footnote 8) and Lemma 17 in the last two steps. ∎
@@
The first inequality follows from the convexity of squaring and the fact that $\left\langle W_{\textnormal{hc},12}\chi|W_{\textnormal{hc},12}\chi\right\rangle\leq\left\langle W_{12}\chi|W_{12}\chi\right\rangle$ pointwise (see Lemma 37 in Appendix A for a proof). The second follows from convexity of $\left\langle W_{12}\chi|W_{12}\chi\right\rangle$ (see Footnote 8 and the proof just referenced) and the boundary condition $|W_{12}(\pm b)\chi|^{2}=b^{2}$.
@@
$\qquad\qquad\quad+E_{F}N(\rho b)^{3}\cdot\rho b^{2}\int\|\tilde{V}\|\Bigg{)}$
@@
$\leq\textnormal{const. }N\rho^{2}N(\rho b)^{4}\left(\left\langle\chi\Big{|}\frac{b}{b-\mathsf{A}^{1,2}}\chi\right\rangle+2b\int\|\tilde{V}\|\right)\leq E_{F}N(\rho b)^{3}\Big{(}\rho b+\rho b^{2}\int\|\tilde{V}\|\Big{)}$.
@@ 5 Lower Bound for scalar potentials
The following lower bound is used to prove Theorem 1. Our method does not allow us to extend to the matrix-valued potentials that we considered for the upper bound—see Footnote 11 for the technical obstruction.
@@
Let $v$ be a radial, repulsive interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) with Neumann boundary conditions. Then, for $\rho R_{0}$, $\rho a_{e}$ and $\rho a_{o}$ sufficiently small, there exists a constant $C_{L}>0$ such that the ground state energy $E_{J}^{\textnormal{Neumann}}(N,L)$ satisfies
@@
where $\epsilon^{\textnormal{LS}}_{J}$ is the minimal energy per site of the spin-$J$ Lai–Sutherland model, i.e. the minimal expectation value of  [^10] We will prove the bound for a chain with open boundary conditions, but this is irrelevant, see Footnote 1.
@@ 5.1.1 The Lieb–Liniger model
$H_{\textnormal{{LL}}}=-\sum_{i=1}^{N}\Delta_{i}+2c\sum_{1\leq i<j\leq N}\delta(x_{i}-x_{j}),$ (5.4)
acting on the Hilbert space $L^{2}_{\text{sym}}([0,L]^{N})$. We write $E_{\textnormal{LL}}^{\textnormal{Neumann}}(N,L,c)$ for its ground state energy at interaction strength $c$ and interval length $L$. In [26, Lemma 17], we proved the following lower bound. The proof uses the exact solution for periodic boundary conditions and a method of Robinson [66] to adapt to Neumann boundary conditions.
@@ 5.2 Proof of the lower bound (Theorem 25)
  • However, this strategy only works for a limited number of particles, and so an argument to extend to any particle number is needed. We use boxes and superadditivity caused by the positive potential.
@@ 5.2.1 Lower bound for small particle numbers
Recall that $v$ is a radial, repulsive interaction compactly supported in $R_{0},R_{0}]$, and that $a_{e}$ and $a_{o}$ are its even and odd wave scattering lengths. For now, let $\Psi$ be some (normalized) wave function describing $N$ fermions (in the domain (2.3) of the energy functional (2.2)), which means that $\Psi(y_{1},\dots,y_{N})$ takes values in $(\mathbb{C}^{2J+1})^{\otimes N}$. We fix a radius $R>R_{0}+a_{o}-a_{e}$ that we will use throughout the proof.
@@
[…]
@@
The purpose of this section is to prove the missing ingredient in the proof of Proposition 30: the bound on the $L^{2}$-norm of $\tilde{\psi}$ in Lemma 29. A proof of this result was given for spinless bosons (or scalar $\Psi$ and $\tilde{\psi}$) in [26, Section 3.2]. We adapt it to the case in which $\Psi$ and $\tilde{\psi}$ take values in $\otimes^{N}\mathbb{C}^{2J+1}$ in this section. To make it easier for the reader to see that the same strategy applies, we restate the lemmas used in [26] and remark on the minor adaptations in the proofs. Note that the only place where we use the assumption that $\Psi$ is the ground state corresponding to $E_{J}^{\textnormal{Neumann}}(N,L)$ is in applying the upper bound Theorem 13 in the last step.
@@
The proof is the same as before, replacing the use of Dyson’s lemma for spinless bosons [26, Lemma 21] by the analogue for spin-$J$ from Lemma 26 (using the bound that only involves $a_{e}$). (Note that the ground state energy of the multicomponent Lieb–Liniger model—for which $\tilde{\psi}$ is a trial state—is simply the bosonic ground state energy. Also note that $R>R_{0}+a_{o}-a_{e}>-a_{e}$ so that $R>|a_{e}|$ by the assumption on $R$ made at the start of Section 5.2.1.)
With Lemma 32, we can prove Lemma 29, which was the missing ingredient in the previous section—we again simply follow our earlier work [26, Proof of Lemma 23]. In the proof, the use of the upper bound [26, Proposition 6] gets replaced the corresponding Theorem 13 (or rather the version for scalar-valued potentials in Corollary 15) and the estimates $a_{e}+(a_{o}-a_{e})\epsilon^{\textnormal{LS}}_{J}\leq a_{o}<R$ and $|a_{e}|<R$.
@@ 5.2.3 Lower bound for arbitrary particle numbers (proof of Theorem 25)
This replaces [26, Corollary 25]. Starting from this corollary and Proposition 30, we now follow the strategy used in [26], where $a$ gets replaced by the infimum from the previous equation. The proof is then completely identical, with the proof of the lower bound Theorem 25 mimicking the proof of [26, Proposition 14], and use of the exact analogue of [26, Lemma 26].

### v2

@@ 4.3.4 Proof of Lemmas 19, 20 and 21
where we used $|W_{12}(x_{1}-x_{2})\chi|^{2}\leq b^{2}$ (see Footnote 9) and Lemma 17 in the last two steps. ∎
@@
The first inequality follows from the convexity of squaring and the fact that $\left\langle W_{\textnormal{hc},12}\chi|W_{\textnormal{hc},12}\chi\right\rangle\leq\left\langle W_{12}\chi|W_{12}\chi\right\rangle$ pointwise (see Lemma 37 in Appendix A for a proof). The second follows from convexity of $\left\langle W_{12}\chi|W_{12}\chi\right\rangle$ (see Footnote 9 and the proof just referenced) and the boundary condition $|W_{12}(\pm b)\chi|^{2}=b^{2}$.
@@
[…]
@@
[…]
@@ 5 Lower Bound for scalar potentials
The following lower bound is used to prove Theorem 1. Our method does not allow us to extend to the matrix-valued potentials that we considered for the upper bound—see Footnote 12 for the technical obstruction.
@@
Let $v$ be an even, positive interaction compactly supported in $R_{0},R_{0}]$, with $a_{e}$ and $a_{o}$ its even and odd wave scattering lengths. Consider a spin-$J$ Fermi gas with energy functional (2.2) and domain (2.3) with Neumann boundary conditions. Then, for $\rho R_{0}$, $\rho a_{e}$ and $\rho a_{o}$ sufficiently small, there exists a constant $C_{L}>0$ such that the ground state energy $E_{J}^{\textnormal{Neumann}}(N,L)$ satisfies
@@
where $\epsilon^{\textnormal{LS}}_{J}$ is the minimal energy per site of the spin-$J$ Lai–Sutherland model, i.e. the minimal expectation value of  [^11] We will prove the bound for a chain with open boundary conditions, but this is irrelevant, see Footnote 2.
@@ 5.1.1 The Lieb–Liniger model
$H_{\textnormal{{LL}}}=-\sum_{i=1}^{N}\partial^{2}_{i}+2c\sum_{1\leq i<j\leq N}\delta(x_{i}-x_{j}),$ (5.4)
acting on the Hilbert space $L^{2}_{\text{sym}}([0,L]^{N})$. We write $E_{\textnormal{LL}}^{\textnormal{Neumann}}(N,L,c)$ for its ground state energy at interaction strength $c$ and interval length $L$. In [25, Lemma 17], we proved the following lower bound. The proof uses the exact solution for periodic boundary conditions and a method of Robinson [78] to adapt to Neumann boundary conditions.
@@ 5.2 Proof of the lower bound (Theorem 25)
  • However, this strategy only works for a limited number of particles, and so an argument to extend to any particle number is needed. We use boxes and superadditivity caused by the positive potential as in [3].
@@ 5.2.1 Lower bound for small particle numbers
Recall that $v$ is an even, repulsive (positive) interaction compactly supported in $R_{0},R_{0}]$, and that $a_{e}$ and $a_{o}$ are its even and odd wave scattering lengths. For now, let $\Psi$ be some (normalized) wave function describing $N$ fermions (in the domain (2.3) of the energy functional (2.2)), which means that $\Psi(y_{1},\dots,y_{N})$ takes values in $(\mathbb{C}^{2J+1})^{\otimes N}$. We fix a radius $R>R_{0}+a_{o}-a_{e}$ that we will use throughout the proof.
@@
After making the $T$-dependent change of coordinates $(y_{1},y_{2},\dots,y_{N})=(x_{1},x_{2}-R_{T_{1}},\dots,x_{N}-\sum^{N-1}_{i=1}R_{T_{i}})$, we find the first inequality in (5.7) as desired. The second inequality follows by noting that the second line of (5.7) is the Lieb–Liniger energy of the state $\psi_{T}$ in a box of length $L-\sum_{i\in T}R_{i}$ with Neumann boundary conditions and coupling $c=2/(R_{0}-a_{o})$, restricted to the particle ordering defined by $M_{T}$. ∎
@@
The purpose of this section is to prove the missing ingredient in the proof of Proposition 30: the bound on the $L^{2}$-norm of $\tilde{\psi}$ in Lemma 29. A proof of this result was given for spinless bosons (or scalar $\Psi$ and $\tilde{\psi}$) in [25, Section 3.2]. We adapt it to the case in which $\Psi$ and $\tilde{\psi}$ take values in $\otimes^{N}\mathbb{C}^{2J+1}$ in this section. To make it easier for the reader to see that the same strategy applies, we restate the lemmas used in [25] and remark on the minor adaptations in the proofs. Note that the only place where we use the assumption that $\Psi$ is the ground state corresponding to $E_{J}^{\textnormal{Neumann}}(N,L)$ is in applying the upper bound Theorem 13 in the last step.
@@
The proof is the same as before, replacing the use of Dyson’s lemma for spinless bosons [25, Lemma 21] by the analogue for spin-$J$ from Lemma 26 (using the bound that only involves $a_{e}$). (Note that the ground state energy of the multicomponent Lieb–Liniger model—for which $\tilde{\psi}$ is a trial state—is simply the bosonic ground state energy. Also note that $R>R_{0}+a_{o}-a_{e}>-a_{e}$ so that $R>|a_{e}|$ by the assumption on $R$ made at the start of Section 5.2.1.)
With Lemma 32, we can prove Lemma 29, which was the missing ingredient in the previous section—we again simply follow our earlier work [25, Proof of Lemma 23]. In the proof, the use of the upper bound [25, Proposition 6] gets replaced the corresponding Theorem 13 (or rather the version for scalar-valued potentials in Corollary 15) and the estimates $a_{e}+(a_{o}-a_{e})\epsilon^{\textnormal{LS}}_{J}\leq a_{o}<R$ and $|a_{e}|<R$.
@@ 5.2.3 Lower bound for arbitrary particle numbers (proof of Theorem 25)
This replaces [25, Corollary 25]. Starting from this corollary and Proposition 30, we now follow the strategy used in [25], where $a$ gets replaced by the infimum from the previous equation. The proof is then completely identical, with the proof of the lower bound Theorem 25 mimicking the proof of [25, Proposition 14], and use of the exact analogue of [25, Lemma 26].